An electric car is being driven through a town at 22 m/s. The kinetic energy of the car is 222640 J. Calculate the mass of the car.
1
Expert's answer
2020-10-05T10:53:39-0400

The kinetic energy

"{\rm KE}=\frac{mv^2}{2}"

Hence, the mass of a car

"m=\frac{2{\rm KE}}{v^2}=\rm\frac{2\times 222640\: J}{(22\: m/s)^2}=920\: kg"